OpenStudy (anonymous):

Find the volume of the cone to the nearest cubic unit. Use a calculator. height 15 cm; radius 18 cm

OpenStudy (campbell_st):

the volume of a cone can be found using the formula \[V = \frac{1}{3}\pi r^2h\] in your question r = 18 and h = 15 just substitute and then evaluate.

OpenStudy (anonymous):

ok thank you!